How to format a submission for DSOM’2001 with the conference’s own L^{A}TEX2e-style

Jens Gustedt^{†} and Somebody Who^{‡} and Some Dummy

INRIA & LORIA, campus scientifique, BP 239, F-54506 Vandœuvre lès Nancy, France gfsjhgfsdjg

Some brief description of your work and how it advances our science.

Keywords: some well classifying words, mandatory!

1 Introduction

The L^{A}TEX2e-style for DSOM’2001 is derived straight forward from the DMTCS style which is also derived from the usual article.sty! Its main purpose is to ensure a common layout policy of all articles in DSOM’2001 and to provide editors, referees and readers with the necessary information. If you think you need an introduction to L^{A}TEX2e or search for pointers to other literature on that, you should consider reference [?] given at the end.

2 Providing Information for the first Page

We need the following type of information:

– The name(s) of the author(s), provided by the \author command.

This is exactly the same as for standard L^{A}TEX2e. Please refer to your L^{A}TEX book to see how this is usually done, or look at the examples given here in this file.

– The title of the manuscript, provided by the \title command.

The title command may be given in two different forms. The first is

\title{Your title goes here}

If done like that, the title that you give is used as running head for the odd numbered pages as well. If your title is too long such that it doesn’t fit into the running head you should use the alternative form

\title[Formatting a submission for DSOM'2001]{How to format a submission
for DSOM'2001 with the journal's own \LaTeX2e-style}

Here the string that inside the [ ] is used in the running head.

– The address(es) of the authors, provided by the \address command.

– Some keywords that classify your work, provided by the \keywords command. Be careful on the choice of these keywords, you are the author, you should know best what is adequate such that your article can be easily and correctly identified by search engines and alike. Give it in the form

\keywords{first item, second, third}

†I am not supported. ‡But he is!

Jens Gustedt and Somebody Who and Some Dummy

So each “key word” might consist of several words in the usual sense. To separate several key words use commas. These keywords must be the same as the ones that are given when you fill out the http-form for submission.

– An abstract of you manuscript, provided by the abstract environment. This should be no longer than a paragraph and concisely reflect the main contributions of your work.

3 Hints for the manuscript itself

3.1 Numbering commands

Please use the standard conventions for all commands and environments that provide a numbering such as theoremlike environments or sections. In particular usual counting starts at 1 and not at 0.

3.2 Proper Names

Please also be careful in the writing of personal names. Customs in different countries are different! Be sure to use a standard transcription of names that use a different alphabet than English, and also be sure to use the full capabilities of L^{A}TEX2e for accentuated character sets that are based on the Latin alphabet. Be sure to catch the correct concept of “last name” in that language.

3.3 Use a Spell Checker

It is considered as being very impolite to leave obvious spelling errors in the manuscript before sending it out. Computers are made for these, use them.

You might either use the North American variant for spelling or the British one, but please don’t mix them in one paper. The same holds for different possible spellings for the same word as for example “acknowledg(e)ment” or “formulae” versus “formulas”. Be coherent.

3.4 Mathematics

– Running text must always constitute correct English phrases.

– All complicated mathematical formulae should be given on separate lines and should not be spread out into the running text. You should use L^{A}TEX2e environments that provide a numbering for such formulae such as equation or eqnarray. Such numbers ease the referee process very much, and after eventual publication easily allow readers to refer to in their own work.

– The quantifiers “∃” and “∀” don’t stand as abbreviations of the partial phrases “there is” and “for all”. They are reserved for logical formulae as such, that is for work that talks itself of logical formulae as a subject.

– The equal sign “=” has different meanings in parts of the two communities that DMTCS addresses.

1. It might stand for mathematical identity that is discovered a posteriori. As an example take the

following phrase:

An easy computation shows that 4! = 24.

2. It might stand for a definition, as in

For convenience, set 0! = 1.

For the later use of “=” Computer Scientist often tend to use “:=”. Referees should be tolerant to these different customs.

4 PDF Files

The style now supports an option pdftex to use in combination with pdflatex. This is only experimental, but we hope that pdflatex will become more stable in the near future. If you are viewing this Formatting a submission for DSOM’2001

document in its pdf form you may see some of the advantages this has: in particular pdf documents produced in that way have included hyperlinks. If you want to know more about these features please refer to

http://xxx.lanl.gov/hypertex/.

If your installation doesn’t support the package hyperref, you should switch of these features by giving the option nohyperref in the \documentclass declaration at the beginning of your manuscript. To give you the possibility to include hyperlinks even if your local installation doesn’t support this, we provide the command \href{URL}{text} in any case.

5 Summary of Options

option description

pdftex to be enabled when processing with pdfla-

tex

notimes switch selection of the times package off

6 Summary of Relevant Commands

command  description

\address the affiliation and address of the authors \keywords a comma separated list of keywords

\qed  produces  □

   \acknowledgements

Références

[OPHS99] Tobias Oetiker, Hubert Partl, Irene Hyna, and Elisabeth Schlegl. The Not So Short Introduction to L^{A}TEX2e, 3.3 edition, 1999. available at

         http://www.loria.fr/services/tex/general/lshort2e.pdf.